A delivery company's charge for an overnight package weighing in excess of one pound given by the formula c=2.83 w+15.51, where w is the weight of package in pounds.
Find the following
A 5 lb package cost ___ to send overnight. Round to the nearest cent.
A 12- ib package cost ___ to send overnight. Round to the nearest cent.

Answer by Mathtut(3670)

c=2.83 w+15.51
:
c(5)=2.83(5)+15.51=14.15+15.51=29.66
:
c(12)=2.83(12)+15.51=33.96+15.51=49.47